HDOffGUI User’s manual

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 Matthias Andree

This document must not be redistributed in any form without prior written consent of the author.

See Chapter 1 [Copying], page 1, for further information.

THERE IS NO WARRANTY FOR THE HDOFFGUI AND HDOFF PACKAGES, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDER OR ANY OTHER PARTY WHICH MAY DIS-TRIBUTE THE PACKAGES PROVIDES THE PROGRAM "AS IS" WITHOUT WAR-RANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PER-FORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DE-FECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL THE COPYRIGHT HOLDER OR ANY OTHER PARTY BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSE-QUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PRO-GRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING REN-DERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF THE COPYRIGHT HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POS-SIBILITY OF SUCH DAMAGES.

The author will however be grateful for bug reports.

Matthias Andree

1 Copying

HDOffGUI 1.4 is copyrighted by Matthias Andree.

Permission to redistribute it is granted if

• the redistributed archive is unchanged^{1} and complete, it is allowed to archive it with: lharc, lha, lzx, zip.

• the package is redistributed without profit, that means: you do not charge more than DM 2,50 (or the equivalent in any other currency) for a disk, DM 30 for a single CD-ROM and DM 60 for a double, triple, quadruple CD-ROM.

• no more than one BBS ad is included, none is allowed on CD-ROM distributions. The only exception is the Tribal BBS in Langenfeld/Rhineland, Germany, whose ad may always be present.

Permission is also granted explicitly to distribute this package on Meeting Pearls, Aminet and Amiga Library Services CD-ROM as well as SaarAG PD disks.

It is strictly prohibited to use this package in commercial packages without licence which can only be obtained from the Author, and the package must not be redistributed by: VOB/Dortmund, Magnamedia/Haar bei München, Pawlowski/Wingst, R2B2/Neukirchen-Vluyn, eagle Computer products, Mallander/Bocholt.

The package MUST NOT be included in cover disks or in disk magazines without prior written consent of the author.

Legal status: Mail-Ware. This means: Although I do not request any monetary charge of you, I will gladly receive donations such as money, registered versions of a shareware tool you have written, postcards.

You must send me an E-Mail, picture postcard or letter if you use HDOffGUI regularly, please state your likes and dislikes and your configuration if you want.

1 this applies also to the file id.diz file!

2 Overview

HDOffGUI is a graphical user interface to HDOff 1.37 or newer. HDOff 1.37 is a hard disk power/noise saver for hard disks connected to the internal A600/A1200/A4000 IDE host. HDOffGUI is written by Gideon Zenz.

HDOffGUI uses triton 1.4 by Stefan Zeiger.

HDOffGUI features:

• Installs itself as Commodity, can be controlled by commodity exchange

• Configuration via tool types and environmental variables

• Can start HDOff asynchronously if it is not already active when HDOffGUI is started • Documentation in .DVI and AmigaGuide format (the master document is written in Texinfo

format and thus can easily be converted to these formats.)

3 Requirements

In order to be able to use HDOffGUI, you need:

• Amiga 600HD/1200HD/4000(T) with at least one IDE hard drive connected to the internal IDE host

• Kickstart/AmigaOS 2.04

• Triton 1.4 by Stefan Zeiger (minimal version^{1} is included)

• HDOff 1.37 by Gideon Zenz (not included!)

You can obtain HDOff 1.37 from either:

• Aminet: ‘disk/misc/HDOff_1.37.lha’
• http://ernie.mi.uni-koeln.de/~gzenz/hdoff.html

1 full version can be obtained from Aminet: ‘util/libs/tri14usr.lha’, a copy is on Aminet 8

3 September 1996 HDOffGUI

4 Installation

There is no Installer script. You do not really need one either, installation is quite easy from Workbench. To install HDOffGUI, follow these steps:

1. If you have PGP and MD5SUM: Check distribution integrity.

• Add my PGP key to your key ring by typing PGP -ka HDOffGUI_34.guide

• Check the signature of ‘MD5.ASC’ by typing PGP MD5.ASC

• Check the files by typing MD5SUM -vc MD5

2. Install HDOff.

3. Install Triton. Double-click on the Triton drawer icon, then double-click the Install icon 4. Drag the HDOffGUI icon anywhere you want, maybe the WBStartup drawer of your boot

partition if you want HDOffGUI to be activated on start-up of your computer. HDOffGUI will automatically start HDOff if it is not active, which means HDOff must reside in C: in order to be found.

5. Configure HDOffGUI. See Section 5.3 [Using], page 7, for details.

5 Configuring and Using HDOffGUI

5.1 Starting HDOffGUI

HDOffGUI can be started both from WorkBench and from Shell. The way it reads its configuration depends on the way it is started, though.

HDOffGUI cannot be started twice. The first copy will pop up its window as soon as you started the second one. The second one leaves immediately.

5.1.1 Starting from Workbench

Double-click the HDOffGUI icon or put HDOffGUI along with its icon in the WBStartup drawer of your boot partition. The configuration is read from tool types.

5.1.2 Starting from Shell

You can just type HDOffGUI and add tool types as command line arguments. Preferably, you would type RUN <>NIL: HDOffGUI, eventually followed by tool types as options on the same line. The tool type options set in the icon are ignored, they must be specified on the command line.

Although the compiler^{1} that I have used allows generation of self-detaching programs (these do not require writing RUN before their names), I have chosen NOT to do so, but instead I recommend starting HDOffGUI from WorkBench^{2}. You could place this line in your ‘s:user-startup’ if you wanted. If you do so, you should not place HDOffGUI in WBStartup, as it would pop up its window.

5.2 Configuring

HDOffGUI can be fully configured via tool types. In addition, the delay time before shutting down the drives can be configured by an environmental variable UNLESS any of TIMEMASTER and TIMESLAVE are set. See HDOFFGUIPREFS for details.

1 SAS/C 6.56 2 Starting a GUI from Shell is somewhat contradictory anyway.

Nota bene: The tool types are only taken into regard if HDOffGUI is started from WorkBench! You need to specify the tool types as arguments if you start HDOffGUI from Shell.

To set Tooltypes, click on the icon and choose ‘Information...’ from the ‘Icons’ menu. You can “comment out” tool types you want to be ignored by enclosing them in parentheses, ‘(’ and ‘)’, example: (TIMESLAVE=10).

You must set at least the SINGLE variable. See below for details.

5.2.1 Configuration via tool types

Attention: the quote marks ‘around this’ are not to be included in any option.

− CX POPKEY specifies the hotkey that pops up the window if it’s hidden. Defaults to

CONTROL LALT H.

− CX POPUP specifies if you want the window shown ‘yes’ or hidden ‘no’ when HDOffGUI

is started. Default: yes (shown). − CX PRIORITY specifies in which order to start commodities. Valid values are -128 to 127,

higher priority commodities are started earlier. − STOPDRIVES specifies the hotkey to stop both drives at once^{3}. Defaults to

CONTROL RALT H.

− STOPMASTER is basically the same as STOPDRIVES, except that it stops the master

drive only (unit 0). − STOPSLAVE is basically the same as STOPDRIVES, except that it stops the slave drive

only (unit 1). This is ignored if SINGLE is set to ‘yes’. − SINGLE specifies if you have one or two drives^{4}. Set to ‘yes’ if you have one drive, set to

‘no’ for two drives. See Section 6.3 [Bugs], page 9, for further information on this. − TIMEMASTER specifies the time^{5} in minutes that the master drive must be unused before

it is switched to standby mode^{6}, set to 0 if you do not want the master drive to be shut down at all. If this tool type is specified, the environmental variable is completely ignored.

− TIMESLAVE is basically the same as TIMEMASTER, but applies to the slave drive. Ignored if SINGLE is ‘yes’.

3 There is actually a short delay between shutting down the master and the slave drive, this is

normal 4 you can of course set this wrong and see what happens. But do not complain HDOffGUI

does not react on Standby or does not show GUI elements for your slave drive :-) 5 IBM says that one spin down/spin up cycle is approximately equivalent to 7 hours of use. 6 The drive will automatically spin up again if accessed

TIMEMASTER and TIMESLAVE are only implemented, because Gideon demanded them. I prefer HDOFFGUI PREFS as the time can be saved to this variable, but not to TIMEMASTER and TIMESLAVE.

5.2.2 Configuring via environmental variable

Nota bene: The environmental variable is ignored if TIMEMASTER or TIMESLAVE is specified. It is also ignored if HDOff is already running when HDOffGUI is started.

The environmental variable is called HDOFFGUI PREFS and must contain two positive, integer numbers, separated by white space, they and have the same meaning as TIMEMASTER and TIMESLAVE (in that order).

Example: saying SETENV HDOFFGUI_PREFS 45 60 before starting HDOff and HDOffGUI would spin down the master drive after 45 minutes being unused, the slave drive after 60 minutes.

The time preferences are saved temporarily^{7} — in ENV: — when you quit the GUI, you can save them permanently using ‘Save preferences’ from the ‘Project’ menu.

5.3 Using the GUI

The GUI window can be (un)zipped by pressing DELETE or by pressing the ZIP gadget (second right in the title bar).

The GUI window can be hidden by Commodoties Exchange, selecting ‘Hide window’ from the ‘Project’ menu, by pressing RCOMMAND H. It can be made appear again by Commodities Exchange or by the CX POPKEY, if specified.

The standby buttons set the corresponding drive to standby mode immediately. They will spin up again automatically if accessed, even if HDOff is no longer active^{8}. The key you can press to take the same action is underlined.

The Time set sliders allow you to change the time the drives need to be unused before they are shut down. A count down is displayed in the ‘Time left’ column. The key you can press to take the same action is the underlined letter from the ‘Drive’ column. If the time is set to

7 to the next reboot 8 General comment: This is different from SCSI drives which can not automatically spin up

again themselves. zero, HDOff will not shut down the drive at all unless you press the Standby button (or its corresponding key).

The other buttons are self-explanatory as are the menu options not explained here. See HDOFFGUI PREFS, for more detailed information on the ‘Save preferences’ menu item.

6 Known Bugs, Restrictions and Incompabitilities

HDOffGUI has been developped and tested on an Amiga 4000 with CyberStorm 68060/50,

24 MB FastRAM, OS 3.0, Conner CFA340A, Western Digital Caviar AC31200F,
GCC 2.7.2^{1}, SAS/C 6.56, GoldEd 3.1, 4.0, 4.1 PRO, Triton 1.4, Triton 2.0 beta, CyberGuard,
MungWall, TexInfo 3.6, PasTeX 1.4.

If you encounter any bugs or misspellings in this document, please contact the author, See Chapter 7 [Author], page 10.

6.1 Known Bugs

• If used together with HDOff 1.36 or earlier versions, these may crash. Use HDOff 1.37 instead.

6.2 Restrictions

6.2.1 Slave recognition

Slave recognition is not yet implemented in HDOff, thus you must configure your drive configuration yourself, see SINGLE. I think this is alright as it needs to be done only upon installation and if you added or removed your slave IDE drive.

6.2.2 Disabling

There is no way to determine if a drive is in standby or running mode if HDOff is disabled for that drive. Thus, the standby button is always available so you can spin down the drive manually at any time.

6.3 Incompatibilites

• If HDEnv is installed, it can cause a drive to spin up again if you quit the GUI or hide it. This is caused by the triton GUI system which saves some parameters when closing a window. I recommend that you use HappyENV instead which is also available from Aminet.

7 Author

If you want to contact the author, be it that you want to send money (cash (only bills), eurocheques and German “Verrechnungsschecks” accepted), bug reports, inquiries, enhancement requests, whatsoever^{1}, here is the address:

Matthias Andree Stormstraße 14 58099 Hagen Federal Republic of Germany

Do not expect replies unless you provide me with a self-addressed, stamped envelope (for Germany) or an International Reply Coupon. E-Mail is the preferred way of contact, but if you send letters/postcards stating your E-Mail address, this will be fine, too.

E-Mail (use either one, but not two or more at the same time):

     HDOffGUI@emma.line.org
     m_andree@tribal.line.org
     mandree@sx1.hrz.uni-dortmund.de

If you are using E-Mail, may I ask you to use PGP if you have installed it. My PGP public key can be found in the AmigaGuide version of this document. Say PGP -ka HDOffGUI_34.guide to extract the PGP key and add it to your key ring.

Before you send bug reports, make sure you have read and understood this document entirely. It may cover certain aspects you have missed so far and solve your problem, thus eliminating the need to contact the author.

If you send bug reports, do state program and triton version, maybe include Segtracker, TNT or Enforcer information, your hardware configuration, what tools, patches and hacks you are using.

1 do not send mail bombs :-)

8 Credit where Credit is due.

Ralph Babel

The Amiga Guru Book (indispensable)

Olaf ‘Olsen’ Barthel

Amiga Developer CD

Ingo ‘Inge’ Dahm

Many AD&D sessions.

Dietmar Eilert

GoldEd PRO

Simon Ferber

He knows why if I say “here’s your pizza”

Oliver Graf

Anubis

Ralf ‘Snoopy’ Imhäuser

Tribal BBS, support

Peter ‘PIK’ Kasel

Many amazing TITAN games.

Kaz a. k. a. Thomas ‘Tiger-Tom’ Kedzior

,,Heute ziehen wir uns richtig einen.”

Marcus ‘Käpt’n’ Kruse

For being a good friend

Kai Posadowsky

many things

Volker Remuß

MetoCloPramiddihydrochlorid? #AmigaGer

Daniel Reymann

total eclipse of the sun

Matthias ‘Schmaußbär’ Schmauß

seeing the midnight sun in Bodø, Norway

Ralph Seichter

no special reason

Christoph ‘Siggi’ Sieg

WET, GET — ,,das ist der Töter”, …

Torsten Spicker

Tribal BBS, special support

Reinhard Spisser

MakeInfo

Volker Stolz MicroDot tools/FAQ, support Gordon ‘Jeppe’ Storkmann

Keep repulsing.

Sascha Tannert

beta testing

Johnny Teveßen

support, Tribal BBS

Manfred Tremmel

nice mails

Sebastiano Vigna

MakeInfo

Oliver Wagner

MicroDot, Voyager

Gideon Zenz HDOff, the idea for HDOffGUI, beta testing

Tribal BBS and its SysAdmins

mail, news, support

DOSIS/2 project and its SysAdmins

mail, news, internet access, home page

SAS Institute

SAS/C Development System 6.51

Free software foundation

GNU tools

Hochschulrechenzentrum der Universität Dortmund

SX1, Internet services

#AmigaGer ‘dictatorial’^{1} amiga celebration channel “do not feed, do not ask about…” :-) on IRC

1 ask L3Xi (Gerrit Hahn) about this Appendix A: Index 13

Appendix A Index

B

Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

C

Configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Configuring, via tool types . . . . . . . . . . . . . . . . . . . . . . . . 6 Copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CX POPKEY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 CX POPUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 CX PRIORITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

D

DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Disabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

E

Environmental variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

G

GUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

H

HDOFFGUI PREFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

I

Incompatibilites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

K

Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

L

Licence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

O

Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

RAMIGA H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 RCOMMAND H. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Restrictions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

S

Shell start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 SINGLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Slave recognition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Starting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Starting, from Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Starting, from WorkBench. . . . . . . . . . . . . . . . . . . . . . . . . 5 STOPDRIVES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 STOPMASTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 STOPSLAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

T

TIMEMASTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 TIMESLAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Tool Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

U

Using the GUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

V

variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

W

Workbench start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Table of Contents

1 Copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Requirements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5 Configuring and Using HDOffGUI. . . . . . . . . . . . . . . . 5

5.1 Starting HDOffGUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.1.1 Starting from Workbench. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.1.2 Starting from Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5.2 Configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.2.1 Configuration via tool types . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5.2.2 Configuring via environmental variable . . . . . . . . . . . . . . . . 7

5.3 Using the GUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6 Known Bugs, Restrictions and Incompabitilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.1 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.2 Restrictions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.2.1 Slave recognition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.2.2 Disabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.3 Incompatibilites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

7 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8 Credit where Credit is due. . . . . . . . . . . . . . . . . . . . . . . 11 Appendix A Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13